### v1

# Bypassing Prompt Injection Detectors through Evasive Injections

Abstract
Large language models (LLMs) are increasingly used in interactive and retrieval-augmented systems, but they remain vulnerable to task drift—deviations from a user’s intended instruction due to injected secondary prompts. Recent work has shown that linear probes trained on activation deltas of LLMs’ hidden layers can effectively detect such drift. In this paper, we evaluate the robustness of these detectors against adversarially optimised suffixes. Using the Greedy Coordinate Gradient (GCG) algorithm, we generate universal suffixes that cause poisoned inputs to evade detection across multiple probes simultaneously. Our experiments on Phi-3 3.8B and Llama-3 8B show that a single suffix can achieve high attack success rates—up to 93.91% and 99.63% respectively when all probes must be fooled, and nearly perfect success (>>99%) under majority-vote criteria. These results demonstrate that activation-delta–based task drift detectors are highly vulnerable to adversarial suffixes, highlighting the need for stronger defences against adaptive attacks. We also propose a defence technique where we generate multiple suffixes and randomly append one of them to the prompts while making forward passes of the LLM and train logistic regression models with these activations. We found this approach to be highly effective against such attacks.
I Introduction
Large language models (LLMs) are increasingly deployed in interactive systems, where they must faithfully follow user instructions. However, LLMs are known to be vulnerable to task drift—deviations from the user’s intended instruction—especially when untrusted text is introduced into their context.
In retrieval augmented generation (RAG) systems [6], [5], LLMs generate their responses based on the data they obtain from external sources. The primary task of the LLM, in this case, is the user’s instruction, which is the only instruction the LLM should execute. However, the data, which the LLM needs to perform the primary task, might contain another instruction for the LLM, which is considered the secondary instruction. Irrespective of whether the secondary instruction is benign or harmful, this is a violation of the concept of data-executable separation. The retrieved data should only provide information for the user’s task; it should never itself be treated as an instruction.
The attacker can stealthily inject a secondary instruction into the data [4], [7], which the LLM might execute, making the system vulnerable and less reliable. This is a significant security risk in RAG systems.
To detect if there is a secondary instruction injected into the retrieved text, Abdelnabi et al. [1] trained logistic regression models on the activation deltas of LLMs’ hidden layers. Their work showed that the difference between the activations before and after processing the text reflects whether the LLM has deviated from the original task.
In this paper, we demonstrate the susceptibility of the linear probes they trained to adversarial attacks. We found that these probes can be fooled into misclassifying poisoned text — text that contains a secondary instruction — as benign by appending an adversarial suffix to the retrieved text. We utilised a well-known algorithm called Greedy Coordinate Gradient (GCG) [10] to find the desired adversarial suffix, which, when appended to the text, can induce the LLM to generate activations that can fool the linear classifiers. Our aim was to find a universal suffix that works on many prompts and for all the internal LLM layers for which classifiers were trained. Additionally, we propose and evaluate two defence mechanisms against such adversarial attacks — PGD-perturbed adversarial training and another approach that utilises adversarial suffixes.
In summary, our work provides the first demonstration that task drift detectors based on activation deltas are vulnerable to adversarial prompt suffixes, highlighting the need for more robust detection mechanisms.
II Background
Prompt Injection Attacks: Prompt injection is a general class of attacks against large language models (LLMs), where an attacker appends or embeds malicious instructions into model inputs [9], [4]. Instead of following the user’s intended task, the LLM may then execute the injected instruction. This undermines the integrity of the system and can lead to unintended or harmful behaviours. Prompt injection can occur in many contexts, ranging from direct user inputs to data retrieved from external sources. Retrieval-augmented generation (RAG) systems are especially vulnerable, since retrieved data can be attacker-controlled, allowing secondary instructions to be stealthily injected into the context. Such injections violate the principle of data–instruction separation, as contextual data should be used only to support the user’s task and never be executed as instructions.
Adversarial Attacks on LLMs: Adversarial examples were first studied in computer vision, where imperceptible perturbations to input images could cause classifiers to mispredict [3]. Extending these ideas to natural language processing (NLP) is more challenging due to the discrete nature of tokens and the requirement to preserve semantic meaning. Nonetheless, a wide range of adversarial strategies have been proposed, including synonym substitution, paraphrasing, character-level perturbations, and suffix-based attacks. In the context of LLMs, adversarial suffixes are especially effective: an attacker can append carefully optimised sequences of tokens to a prompt to manipulate the model’s behaviour or evade detection. One algorithm for generating such suffixes is the Greedy Coordinate Gradient (GCG) method [10], which iteratively updates suffix tokens by estimating gradients and selecting the most promising replacements.
Task Drift Detection with Linear Probes: To defend against prompt injection, Abdelnabi et al. [1] introduced the concept of task drift detection. Their method attaches logistic regression classifiers (linear probes) to multiple hidden layers of an LLM, trained on the activation deltas between clean and poisoned inputs. The intuition is that if a model begins to follow a secondary instruction, the internal activations diverge from those induced by the user’s primary instruction. By monitoring these activation deltas, the probes can distinguish clean prompts from poisoned ones with high accuracy. This approach provides a lightweight and model-agnostic detection mechanism, as the probes are simple linear classifiers operating on fixed representations.
Our work builds directly on this framework. While Abdelnabi et al. [1] evaluated robustness against straightforward prompt injection, we investigate whether the same linear probes can be circumvented through adversarial optimisation of suffixes.
III Threat Model
We assume a general-purpose retrieval-augmented language model (LLM), similar to the one described in [1]. The user can instruct the LLM to perform a task — the primary instruction. The LLM requires and obtains external data to successfully execute user’s instruction. This data can be classified into two categories: clean and poisoned. Clean data contains only factual or contextual information relevant to the user’s task. Poisoned data, in contrast, contains an injected secondary instruction intended to influence the LLM. This injected instruction can be either malicious (e.g., “how to make a bomb?”) or benign (e.g., “what is the White House?”). Even a benign instruction is unwanted in this context because of the violation of data-instructions separation.
For example, suppose the user asks in 2025, “Who is the president of the USA?” If the LLM was trained before the 2024 U.S. election, it will be unaware of the change in presidency and must rely on retrieved data. If the retrieved passage contains an additional instruction such as “What is the White House?”, the model is induced to process an unintended instruction. Even though this secondary instruction is benign, it still constitutes task drift, as the model deviates from the user’s original intent.
Because the retrieved data originates from external, potentially untrusted sources, we assume the attacker controls the data channel. The attacker can therefore inject secondary instructions into the retrieved text. However, the attacker cannot modify the user’s primary instruction, nor can they alter the parameters of the LLM or the task drift detectors.
We distinguish between two attacker models:
  1. 1.
    Baseline attacker (as in Abdelnabi et al. [1]): This attacker injects a secondary instruction into the retrieved text, with the goal of making the LLM execute it. This is the threat model considered in the original task drift detection work.
  2. 2.
    Adaptive attacker (our extended formulation): We consider a stronger attacker who, in addition to injecting a secondary instruction, appends an adversarial suffix to the poisoned text. This suffix is optimised to fool the linear classifiers trained to detect task drift. By bypassing detection, the poisoned instruction has a higher chance of being executed by the LLM without raising alarms.
The defender deploys linear probes trained on LLM activation deltas at specific layers. The defender’s goal is to detect task drift whenever a secondary instruction is present in the retrieved data.
We define attack success as follows: a poisoned input is classified as benign by all (or a majority of) the drift detectors, despite containing a secondary instruction. In the baseline setting, the attacker’s success depends only on making the LLM execute the injected instruction. In the adaptive setting, success requires simultaneously evading detection and preserving the effectiveness of the injected instruction.
IV Victim Models
There are multiple task drift detectors trained on the activations of different internal layers of a specific LLM. For example, there are 5 classifiers for Phi-3 3.8B and Llama-3 8B trained on layers 0, 7, 15, 23, and 31. While the first classifier performs poorly and misclassifies many prompts, others detect task drift with very high accuracy. Table I shows the accuracy of the linear probes on the test clean and test poisoned datasets.
Layer Phi-3 3.8B Llama-3 8B
Clean Poisoned Clean Poisoned
0 71.56% 57.31% 66.99% 64.79%
7 97.39% 98.54% 92.83% 83.53%
15 78.45% 99.46% 92.54% 96.67%
23 80.45% 99.91% 94.89% 97.94%
31 93.06% 99.37% 91.29% 94.59%
TABLE I: Linear models’ accuracy on test clean and test poisoned dataset
Same Prediction Phi-3 3.8B Llama-3 8B
Clean Poisoned Clean Poisoned
All five classifiers 48.27% 56.54% 60.84% 59.76%
Last four classifiers 68.26% 97.80% 83.41% 79.38%
TABLE II: Same prediction by linear models of Phi-3 and Llama-3 for test clean and test poisoned dataset
Table II shows the consistency of the linear models in their predictions. For example, 97.80% for the last four classifiers of Phi-3 on the poisoned dataset means that for 97.80% of the poisoned test prompts, all four classifiers agreed on the same label — either 1 (task drift) or 0 (no drift). It, however, does not mean that all these predictions are correct. It only means that the classifiers predicted the same label, regardless of whether that prediction is correct or not.
Since the first classifier makes a lot of mistakes and its predictions are not consistent with the others, it can reasonably be ignored while defending against prompt injections. However, the other classifiers detect task drift with high accuracy and consistency, and the defender can rely on either the same prediction by all four classifiers or a majority vote among them. Therefore, the attacker has to fool all or at least most of the classifiers to successfully perform prompt injection attacks. Otherwise, even if one classifier is fooled, the others will still detect task drift and the defender will foil the attack.
V Proposed Approach
V-A Attack
In this paper, we propose an adversarial attack aimed at fooling all the classifiers simultaneously. We want to append a suffix to the prompt that contains a secondary instruction, and make all the classifiers classify it as benign. We treat this as a classification problem, where the goal is to drift the output of the classifier away from the true class label 1 to the target class label 0. Since we optimise the suffix for all the classifiers simultaneously, we therefore add the gradients of the losses of all the classifiers at their respective hidden layers with the gradients that come from the next hidden layer during backpropagation, and update the tokens of the suffix based on these accumulated gradients, as illustrated in Fig. 1.
Figure 1: Gradient Computation
Formally, the gradient of the loss with respect to activations at layer $l$ has two components: (i) the gradient propagated from the next layer, and (ii) an additional gradient from the linear classifier attached at layer $l$.
$\boxed{\frac{\partial\mathcal{L}}{\partial h^{(l)}}=\underbrace{\left(\frac{\partial h^{(l+1)}}{\partial h^{(l)}}\right)^{\top}\frac{\partial\mathcal{L}}{\partial h^{(l+1)}}}_{\text{from next layer}}+\underbrace{\mathbf{1}_{\{l\in\mathcal{P}\}}\cdot(\sigma(z^{(l)})-y)W^{(l)}}_{\text{from classifier at layer $l$}}}$ (1)
Thus, our proposed attack jointly optimises the adversarial suffix across all probes, forcing consistency in misclassification.
V-B Defence
We experimented with two approaches to defend against such suffix-based adversarial attacks. The first one is the standard PGD-perturbed adversarial training in which we trained logistic regression models using PGD-perturbed poisoned activations along with the clean and poisoned activations. Since these models have not seen any suffix during their training because no suffix has been used to train them, therefore any optimised suffix that can fool the baseline classifiers can be used to test their robustness against suffix attacks.
The second approach involves training the linear models with the activations obtained by appending suffixes to the prompts. Firstly, we started with different initial suffixes and generated multiple suffixes by targeting the baseline models using the algorithm described in V-A and split them into train, validation, and test set. Then, we randomly appended one of the train and validation suffixes to the training and validation prompts respectively and made a forward pass of the LLM to generate activations, which we refer to as adversarially poisoned activations. Finally, we trained logistic regression models with these activations along with the clean and poisoned activations. We tested the robustness of these models by targeting them with test suffixes that the models have not seen before. Furthermore, we also tried to optimise suffixes by directly attacking the adversarially trained models. However, we found that these models are extremely resistant to this type of attack. It’s very difficult to find a suffix despite starting from different initial suffixes and attacking the models directly.
VI Experiments
VI-A Setup
We conducted our experiments on two LLMs and their associated linear probes: Phi-3 3.8B [2] and Llama-3 8B [8]. The probes were logistic regression classifiers trained on activation deltas at layers 0, 7, 15, 23, and 31, as reported in [1].
We recreated the TaskTracker dataset used in [1], which consists of clean prompts and poisoned prompts containing secondary instructions. We followed the same train/test splits as in [1]. There are 418,110 prompts in the train subset and 31,134 prompts in the test poisoned subset.
For the attack, we used the Greedy Coordinate Gradient (GCG) algorithm to optimise an adversarial suffix. We restricted the suffix length to the length of the initial suffix. We also tried allowing the length of the suffix of Llama to vary from the length of the initial suffix, and it performed equally well. However, after a few iterations, the suffix tended to become excessively long, so we fixed the suffix length across iterations for consistency.
We began with the first prompt from the list of randomly chosen training prompts and optimised the suffix for this prompt. We added the next prompt after finding a suffix that could make all the classifiers predict the wrong label for the current prompt with a certain confidence threshold (e.g., 0.7). We repeated this process, and if the classifiers misclassified 80% or more prompts after an iteration, the next prompt was added. Even though we intended to run the algorithm for 500 iterations, we found that it was usually possible to obtain a suffix that worked on many prompts after at most 100–120 iterations.
There are multiple ways to evaluate attack success. We considered the following three criteria:
  1. 1.
    Fooling all the probes simultaneously
  2. 2.
    Fooling the majority of the probes
  3. 3.
    Fooling last four probes, since the first probe is unreliable and often makes errors even without an adversarial suffix
Defence experiment was conducted on Phi-3 only and the performance is evaluated by how correctly the adversarially trained models detect the presence of secondary instructions when an adversarial suffix is appended to the prompts along with the correct classification of clean and poisoned (without suffix) prompts. We consider the following three criteria to evaluate defence success:
  1. 1.
    Correct classification by all the probes
  2. 2.
    Correct classification by the majority of the probes
  3. 3.
    Correct classification by the last four probes
VI-B Results
We evaluate the effectiveness of the proposed adversarial suffix against the task drift detectors for both Phi-3 3.8B and Llama-3 8B. Results are reported in terms of the number of classifiers fooled for each prompt, as well as layer-wise misclassification rates.
VI-B1 Attack Success Rate
Table I reports the attack success rate (ASR) under different evaluation metrics. For Phi-3, the suffix achieves an ASR of 93.91% when requiring all five probes to be fooled simultaneously, and 95.28% when considering the last four probes only. Under a majority-vote criterion ($\geq$3 probes), the ASR rises to 99.79%. For Llama-3, the attack is even more effective, reaching 99.63% ASR for both “all five” and “last four” metrics, and nearly perfect success (99.99%) under the majority-vote criterion. These results show that adversarial suffixes can reliably evade probe-based detectors.
Success metric Phi-3 3.8B (%) Llama-3 8B (%)
All five probes 93.91 99.63
Last four probes 95.28 99.63
Majority ($\geq$3) 99.79 99.99
TABLE III: Attack Success Rate (ASR) across evaluation metrics
VI-B2 Distribution of Misclassifications
Fig. 2 shows the distribution of prompts misclassified by different numbers of classifiers for Phi-3, before and after the attack. Without suffixes, most poisoned prompts were correctly detected, with 56.56% flagged by all probes and 41.83% flagged by four probes. However, with the adversarial suffix, 93.91% of prompts were misclassified by all probes simultaneously, demonstrating the effectiveness of the attack.
Figure 2: Total number of prompts misclassified by a specific number of classifier Phi-3 3.8B
A similar pattern holds for Llama-3 (Fig. 3). Without suffixes, only 0.17% of prompts fooled all probes, while 59.59% were correctly detected by all probes. After applying the suffix, 99.63% of prompts successfully evaded detection by all probes.
Figure 3: Total number of prompts misclassified by a specific number of classifier Llama-3 8B
VI-B3 Layer-wise Misclassification
Fig. 4 and 5 show layer-wise results. For Phi-3, probes at deeper layers were initially very accurate, with misclassification rates below 1% on poisoned prompts at layers 7, 15, 23, and 31. After the attack, these rates increased dramatically, exceeding 95% for every deep layer. The Llama-3 detectors were similarly robust in the clean setting, with error rates below 6% across all but the first layer. With the adversarial suffix, however, nearly every probe was fooled, with misclassification rates of 99–100% across layers.
Figure 4: Layer wise misclassification Phi-3 3.8B
Figure 5: Layer-wise misclassification Llama-3 8B
These results demonstrate that the activation-delta–based linear probes, though effective in detecting straightforward prompt injections, are highly vulnerable to adaptive adversarial attacks. A single optimised suffix can simultaneously mislead multiple probes across different layers, achieving near-universal evasion on both Phi-3 and Llama-3 models.
VI-C Defence Performance
Table IV and V show the layer-wise performance on clean and poisoned dataset respectively. For clean dataset, suffix trained models’ performance was on par with the baseline models except for the first layer’s classifier whose performance was really poor, while the PGD-trained models performed much worse. For poisoned dataset, PGD-trained models performed best, with the first two classifiers achieving perfect accuracy, while the suffix-trained models’ performance was very close to that of baseline models.
Layer Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
0 71.56% 0% 0.16%
7 97.39% 0% 91.77%
15 78.45% 57.53% 88.14%
23 80.45% 59.47% 84.40%
31 93.06% 29.21% 82.31%
TABLE IV: Linear Models’ Layer-wise Performance on Test Clean Dataset — Phi-3 3.8B
Layer Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
0 57.31% 100% 99.37%
7 98.54% 100% 97.85%
15 99.46% 99.38% 98.26%
23 99.91% 99.88% 99.58%
31 99.37% 99.59% 97.82%
TABLE V: Linear Models’ Layer-wise Performance on Test Poisoned (without suffix) Dataset — Phi-3 3.8B
Table VI and VII show the performance of linear models across different evaluation metrics for clean and poisoned dataset respectively. For the strictest metric, where all probes must classify correctly, adversarially trained models perform poorly on clean dataset because of the poor performance of the first classifier, while they outperform the baseline models on poisoned dataset.
Metric Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
All probes 47.92% 0% 0.13%
Last four probes 67.77% 0% 73.04%
Majority ($\geq$3) 94.32% 18.42% 85.25%
TABLE VI: Linear Models’ Performance across evaluation metrics on Test Clean Dataset — Phi-3 3.8B
Metric Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
All probes 56.53% 99.04% 94.94%
Last four probes 97.78% 99.04% 95.49%
Majority ($\geq$3) 99.74% 99.98% 99.47%
TABLE VII: Linear Models’ Performance across evaluation metrics on Test Poisoned (without suffix) Dataset — Phi-3 3.8B
Table VIII, IX, and X show how different types of linear models fared in defending against adversarial attacks where test suffixes were appended to the prompts. While the baseline models invariably failed in every metric, PGD-trained models showed varying level of performance depending on the suffix and metric. Suffix-trained models demonstrated satisfactory and reliable performance, with accuracy of 80-100% based on suffix and metric considered.
Suffix Index Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
1 0.00% 36.67% 79.79%
2 0.00% 38.09% 86.73%
3 0.02% 47.61% 98.36%
4 0.02% 17.77% 99.41%
5 0.00% 49.49% 94.95%
6 0.00% 1.31% 99.38%
7 0.01% 42.57% 97.33%
8 0.09% 44.63% 93.53%
9 0% 11.82% 83.03%
10 0.17% 41.93% 81.57%
TABLE VIII: Linear models’ performance in defending against adversarial attacks with test suffixes (All five probes)
Suffix Index Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
1 0.02% 36.67% 79.79%
2 0% 38.09% 86.73%
3 0.03% 47.61% 98.36%
4 0.04% 17.77% 99.41%
5 0.01% 49.49% 94.95%
6 0.01% 1.31% 99.38%
7 0.01% 42.57% 97.33%
8 0.19% 44.63% 93.53%
9 0.01% 11.82% 83.03%
10 0.18% 41.93% 81.57%
TABLE IX: Linear models’ performance in defending against adversarial attacks with test suffixes (Last four probes)
Suffix Index Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
1 0.23% 83.98% 100.00%
2 0.40% 58.01% 98.81%
3 0.24% 83.37% 100.00%
4 0.40% 49.10% 100.00%
5 0.41% 94.38% 100.00%
6 1.60% 52.89% 100.00%
7 0.30% 65.32% 100.00%
8 1.37% 65.51% 99.99%
9 0.11% 67.07% 99.59%
10 0.57% 79.69% 99.87%
TABLE X: Linear models’ performance in defending against adversarial attacks with test suffixes (Three or more probes)
The extremely poor performance of PGD-trained models on clean dataset and much worse performance in defending against suffix attacks than suffix-trained models is possibly because of the differences between the actual activations obtained by appending a suffix to the prompts and making forward passes of the LLM and the activations obtained by adversarially perturbing the poisoned activations using PGD. The latter might differ a lot from the real ones, which contributes to the ineffectiveness of PGD-trained models and they perform poorly when they classify real activations.
VII Conclusion
In this work, we investigated the robustness of task drift detection using activation-delta–based linear probes. While these probes were shown to be highly effective in detecting straightforward prompt injection attacks, our experiments demonstrate that they are vulnerable to adaptive adversarial attacks. By optimising a universal suffix using the Greedy Coordinate Gradient algorithm, we were able to induce high misclassification rates across all probes and layers, effectively bypassing detection in both Phi-3 3.8B and Llama-3 8B. These findings highlight a significant limitation in current approaches to task drift detection and underscore the need for more resilient defences.
Additionally, we introduced a defence technique to defend against adversarial attacks that utilise suffixes. Our approach proved to be highly effective. The linear models we have trained with adversarial activations demonstrated strong resistance against suffix attacks, not only in detecting task drift when test suffixes were appended to the prompts, but also when we tried to optimise suffixes by attacking these models directly.